Grade 11Trigonometry

plz help me to solv this. if A+B+C=pie ,cosA=cosBcosC , prove that tanA=tanB+tanC

sudipta biswas
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
R.H.S.
= tanB+tanC
= (sinB/cosB)+(sinC/cosC)
= (sinBcosC+cosBsinC)/(cosB.cosC)     
= sin(B+C)/cosA            [cosA=cosB.cosC]
= sinA/cosA                    [A+B+C=pie, B+C=pie-A, sin(B+C)=sin(pie-A)= sinA]
= tanA = L.H.S. Proved